• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Run coordinator

    Log entry time 02:15:21 on February16,2002

    Entry number 82872

    keyword=Status and upcoming shifts

    The target rapair started at about 00:00 on Saturday, Feb 16.
    The plan is to swap loops. By Saturday late evening the target
    should be cooled down and could be tested again.

    The shifts' status:
    Saturday:
    owl,day - cancelled
    swing - standby
    Sunday:
    owl - standby

    More information is expected by the end of Saturday.